# Structure-Preserving Oscillation-Eliminating Discontinuous Galerkin Schemes for Ideal MHD Equations: Locally Divergence-Free and Positivity-Preserving

### v1

@@ 2.2 LDF DG finite element spaces
where $\boldsymbol{x}=\left(x,y\right)$ denotes the spatial coordinate vector, and $\mathbb{P}^{k}(I_{ij})$ is the space of polynomials of total degree less than or equal to $k$ on the element $I_{ij}$. An orthogonal basis of $\mathbb{P}^{k}(I_{ij})$ is denoted as $\{\phi_{ij}^{(\boldsymbol{\alpha})}(\boldsymbol{x}):|\boldsymbol{\alpha}|\leq k\}$, where $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2})$ is the multi-index vector with $|\boldsymbol{\alpha}|=\alpha_{1}+\alpha_{2}$. For example, we adopt the following (orthogonal) Legendre basis over cell $I_{ij}$:
@@
$\left(0,0\right),\quad\left(1,0\right),\quad\left(0,1\right),\quad\left(2,0\right),\quad\left(1,1\right),\quad\left(0,2\right),\quad...,$
@@
We seek the divergence-free approximation to $\boldsymbol{Q}$ in the following LDF space [7]:
@@
The local orthogonal basis of $\mathbb{V}_{Q}^{k}$ on element $I_{ij}$ is denoted as $\left\{\boldsymbol{\varphi}_{ij}^{\left(\mu\right)}\left(\boldsymbol{x}\right),\mu=0,1,\cdots,D_{Q}^{k}-1\right\}$, where $D_{Q}^{k}=\left(k+1\right)\left(k+4\right)/2$ is the dimensionality of the LDF polynomial space defined on $I_{ij}$. We employ the following local orthogonal LDF basis
@@ 2.3 Semi-discrete LDF DG formulation
where the superscripts -- and $+$ indicate the left- and right-hand-side limits at the cell interfaces, and
@@
where $\left\{x_{i}^{(\mu)}\right\}_{\mu=1}^{q}$ and $\left\{y_{j}^{(\mu)}\right\}_{\mu=1}^{q}$ are the $q$-point Gauss quadrature nodes with $2q-1\geq 2k$ in $\left[x_{i-\frac{1}{2}},x_{i+\frac{1}{2}}\right]$ and $\left[y_{j-\frac{1}{2}},y_{j+\frac{1}{2}}\right]$, respectively, and $\left\{\omega_{\mu}^{G}\right\}_{\mu=1}^{q}$ are the associated weights.
@@
$-\int_{x_{i-\frac{1}{2}}}^{x_{i+\frac{1}{2}}}\left(\boldsymbol{v}\left(x,y_{j+\frac{1}{2}}^{-}\right)\cdot\widehat{\boldsymbol{F}_{2}^{Q}}_{,j+\frac{1}{2}}\left(x,t\right)-\boldsymbol{v}\left(x,y_{j-\frac{1}{2}}^{+}\right)\cdot\widehat{\boldsymbol{F}_{2}^{Q}}_{,j-\frac{1}{2}}\left(x,t\right)\right)\mathrm{~{}d}x$
@@
$\widehat{\boldsymbol{F}_{2}^{Q}}_{,j+\frac{1}{2}}\left(x,t\right):=\widehat{\boldsymbol{F}_{2}^{Q}}\left(\boldsymbol{U}_{h}\left(x,y_{j+\frac{1}{2}}^{-},t\right),\boldsymbol{U}_{h}\left(x,y_{j+\frac{1}{2}}^{+},t\right)\right)\$
@@ 2.4 Fully discrete OEDG schemes
The combination of semi-discrete schemes (9) and (12) can be regarded as an ordinary differential equation (ODE) with respect to time $t$. For brevity, we rewrite equations (9) and (12) in the following ODE form:
@@
where $\boldsymbol{U}_{h}$ denotes the combination of $\boldsymbol{R}_{h}$ and $\boldsymbol{Q}_{h}$. This ODE system (13) is typically discretized in time by a Runge-Kutta method, for example, the explicit third-order strong-stability-preserving (SSP) Runge-Kutta method:
@@
where $\boldsymbol{U}_{\sigma}^{n}$ is the modified numerical solution at the $n$-th time level, and we define $\boldsymbol{U}_{\sigma}^{0}=\boldsymbol{U}_{h}^{0}$.
@@ 2.4.1 Component-wise OE procedure for $\boldsymbol{R}_{h}$
where $\tau$ is a pseudo-time different from $t$, ``"``"`$\circ$"` ` ∘ " denotes the Hadamard product of two vectors, $P^{m}$ is the standard $L^{2}$-projection operator into $\mathbb{V}_{R}^{m}$ for $m\geq 0$, and we define $P^{-1}=P^{0}$. The vector $\boldsymbol{\delta}_{ij}^{m}\left(\boldsymbol{R}_{h}^{n,\ell}\right)$ is defined as
@@
where $\beta_{ij}^{x}$ and $\beta_{ij}^{y}$ denote suitable estimates of the local maximum wave speed in the $x$- and $y$-directions, respectively, and $\boldsymbol{\sigma}_{i+\frac{1}{2},j}^{m}\left(\boldsymbol{R}_{h}\right)$ and $\boldsymbol{\sigma}_{i,j+\frac{1}{2}}^{m}\left(\boldsymbol{R}_{h}\right)$ correspond to the damping coefficients on the $x=x_{i+\frac{1}{2}}$ and $y=y_{j+\frac{1}{2}}$ edges, respectively. In practice, we take $\beta_{ij}^{x}$ and $\beta_{ij}^{y}$ as the spectral radii of the Jacobian matrices $\frac{\partial\boldsymbol{F}_{1}}{\partial{\boldsymbol{U}}}\left(\bar{\boldsymbol{U}}_{h}^{n,\ell}\right)$ and $\frac{\partial\boldsymbol{F}_{2}}{\partial{\boldsymbol{U}}}\left(\bar{\boldsymbol{U}}_{h}^{n,\ell}\right)$, respectively, and $\bar{\boldsymbol{U}}_{h}^{n,\ell}$ denotes the cell average of $\boldsymbol{U}_{h}^{n,\ell}\left(\boldsymbol{x}\right)$ over $I_{ij}$. In equation (17), the damping coefficients are given by
$\boldsymbol{\sigma}_{i+\frac{1}{2},j}^{m}\left(\boldsymbol{R}_{h}\right)=\left(\sigma_{i+\frac{1}{2},j}^{m}\left(R_{h}^{\left(1\right)}\right),\sigma_{i+\frac{1}{2},j}^{m}\left(R_{h}^{\left(2\right)}\right),\cdots,\sigma_{i+\frac{1}{2},j}^{m}\left(R_{h}^{\left(N_{R}\right)}\right)\right)^{\top}$
and
$\boldsymbol{\sigma}_{i,j+\frac{1}{2}}^{m}\left(\boldsymbol{R}_{h}\right)=\left(\sigma_{i,j+\frac{1}{2}}^{m}\left(R_{h}^{\left(1\right)}\right),\sigma_{i,j+\frac{1}{2}}^{m}\left(R_{h}^{\left(2\right)}\right),\cdots,\sigma_{i,j+\frac{1}{2}}^{m}\left(R_{h}^{\left(N_{R}\right)}\right)\right)^{\top},$
@@
Note that the coefficient $\boldsymbol{\delta}_{ij}^{m}\left(\boldsymbol{R}_{h}^{n,\ell}\right)$ in the damping equations 16 only depends on the “initial” value $\boldsymbol{R}_{\sigma}\left(\boldsymbol{x},0\right)=\boldsymbol{R}_{h}^{n,\ell}\left(\boldsymbol{x}\right)$. Consequently, the damping equations 16 are actually a linear ODE system, and its exact solution can be analytically given. We now derive the exact solver of damping equations 16 for $\boldsymbol{R}_{\sigma}\left(\boldsymbol{x},\tau\right)$. Based on the local orthogonal basis $\left\{\phi_{ij}^{\left(\boldsymbol{\alpha}\right)}\left(\boldsymbol{x}\right):\left|\boldsymbol{\alpha}\right|\leq k\right\}$ of $\mathbb{P}^{k}\left(I_{ij}\right)$, the “initial” value $\boldsymbol{R}_{h}^{n,\ell}\left(\boldsymbol{x}\right)$ of the damping equations (16) can be expanded as
@@
Substitute (20)-(21) into (16) and take $v=\phi_{ij}^{\left(\boldsymbol{\alpha}\right)}\left(\boldsymbol{x}\right)$. For $\left|\boldsymbol{\alpha}\right|=\mu\geq 1$, we get
@@
Solving (22) gives
@@
$\int_{I_{ij}}\left(\boldsymbol{R}_{\sigma}-P^{-1}\boldsymbol{R}_{\sigma}\right)\phi_{ij}^{\left(\boldsymbol{\alpha}\right)}\left(\boldsymbol{x}\right)\mathrm{~{}d}\boldsymbol{x}=\int_{I_{ij}}\left(\boldsymbol{R}_{\sigma}-P^{0}\boldsymbol{R}_{\sigma}\right)\phi_{ij}^{\left(\boldsymbol{\alpha}\right)}\left(\boldsymbol{x}\right)\mathrm{~{}d}\boldsymbol{x}\equiv\boldsymbol{0},$
we have
$\frac{\mathrm{d}}{\mathrm{d}\tau}\hat{\boldsymbol{R}}_{ij}^{\left(\boldsymbol{\alpha}\right)}\left(\tau\right)\int_{I_{ij}}\left(\phi_{ij}^{\left(\boldsymbol{\alpha}\right)}\left(\boldsymbol{x}\right)\right)^{2}\mathrm{~{}d}\boldsymbol{x}=\boldsymbol{0},$
@@
The OE procedure (25) implemented after each Runge-Kutta stage modifies the high-order moments of numerical polynomials. Because this OE procedure does not interfere with DG spatial discretization and RK stage update, it can be easily incorporated to existing DG codes as an independent module. Only simple addition and multiplication operations for cell averages and derivatives are involved, allowing this OE procedure to be extended to general meshes straightforwardly and efficiently. Note that the OE procedure is applied directly to the conservative variables without the need for characteristic decomposition, which greatly simplifies the implementation and reduces the computational costs. It is observed that the OE procedure contains information about wave speed and scales linearly with respect to the solution magnitude, resulting in being capable of eliminating spurious oscillations effectively for problems across different scales and wave speeds. Furthermore, as described in [1], the OE method remains stable under the normal CFL condition, and many desirable properties of the original DG method are preserved, including conservation, high-order accuracy, optimal fully-discrete error estimates in theory, and numerical superconvergence for linear scalar conservation laws.
@@ 2.4.2 LDF OE procedure for $\boldsymbol{Q}_{h}$
$\begin{split}\boldsymbol{Q}_{\sigma}^{n,\ell}\left(\boldsymbol{x}\right)&=\left(\mathcal{F}_{\text{OE}}\boldsymbol{Q}_{h}^{n,\ell}\right)\left(\boldsymbol{x}\right)=\boldsymbol{Q}_{\sigma}\left(\boldsymbol{x},\Delta t_{n}\right)\\ &=\sum_{\mu=0}^{D_{Q}^{0}-1}Q_{ij}^{\left(\eta\right)}\left(0\right)\boldsymbol{\varphi}_{ij}^{\left(\eta\right)}\left(\boldsymbol{x}\right)+\sum_{\mu=1}^{k}\exp\left(-\Delta t_{n}\sum_{m=0}^{\mu}\delta_{ij}^{m}\left(\boldsymbol{Q}_{h}^{n,\ell}\right)\right)\sum_{\eta=D_{Q}^{\mu-1}}^{D_{Q}^{\mu}-1}Q_{ij}^{\left(\eta\right)}\left(0\right)\boldsymbol{\varphi}_{ij}^{\left(\eta\right)}\left(\boldsymbol{x}\right).\end{split}$ (35)
The LDF OE procedure modifies the high-order moments of original LDF magnetic field. Thanks to the exact divergence-free solution of OE solver in the LDF space, the modified magnetic field naturally retains the LDF property of the original magnetic field. Moreover, this LDF OE procedure also inherits many notable advantages mentioned in the component-wise OE procedure, including non-intrusive to be absorbed into existing LDF DG codes, straightforward and efficient to be extend on general meshes, stable under normal CFL condition, scale-invariant and evolution-invariant in eliminating the spurious oscillations for problems across various scales and wave speeds, free of characteristic decomposition, and retaining desirable conservation, high-order accuracy, and optimal convergence rates of the original DG method. All these advantages highlight the strengths of LDF OE procedure again.

### v2

@@ 2.2 LDF DG finite element spaces
where $\boldsymbol{x}=\left(x,y\right)$ denotes the spatial coordinate vector, and $\mathbb{P}^{k}(I_{ij})$ is the space of polynomials of total degree less than or equal to $k$ on the element $I_{ij}$. An orthogonal basis of $\mathbb{P}^{k}(I_{ij})$ is denoted as $\{\phi_{ij}^{(\boldsymbol{\alpha})}(\boldsymbol{x}):|\boldsymbol{\alpha}|\leq k\}$, where $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2})$ is the multi-index vector with $|\boldsymbol{\alpha}|=\alpha_{1}+\alpha_{2}$. For example, we adopt the following local orthogonal Legendre basis over cell $I_{ij}$:
@@
$\left(0,0\right),\quad\left(1,0\right),\quad\left(0,1\right),\quad\left(2,0\right),\quad\left(1,1\right),\quad\left(0,2\right),\quad\dots$
@@
We seek a divergence-free approximation to $\boldsymbol{Q}$ in the following LDF space [6]:
@@
The local orthogonal basis of $\mathbb{V}_{Q}^{k}$ on element $I_{ij}$ is denoted as $\{\boldsymbol{\varphi}_{ij}^{\left(\mu\right)}\left(\boldsymbol{x}\right),\mu=0,1,\cdots,D_{Q}^{k}-1\}$, where $D_{Q}^{k}=\left(k+1\right)\left(k+4\right)/2$ is the dimensionality of the LDF polynomial space defined on $I_{ij}$. We employ the following local orthogonal LDF basis
@@ 2.3 Semi-discrete LDF DG formulation
where the superscripts “--” and “$+$” indicate the left- and right-hand-side limits at the cell interfaces, respectively, and
@@
where $\left\{x_{i}^{(\mu)}\right\}_{\mu=1}^{q}$ and $\left\{y_{j}^{(\mu)}\right\}_{\mu=1}^{q}$ are the $q$-point Gauss quadrature nodes with $2q-1\geq 2k$ in $\left[x_{i-\frac{1}{2}},x_{i+\frac{1}{2}}\right]$ and $\left[y_{j-\frac{1}{2}},y_{j+\frac{1}{2}}\right]$, respectively, and $\left\{\omega_{\mu}^{G}\right\}_{\mu=1}^{q}$ are the associated weights for the interval $[-\frac{1}{2},\frac{1}{2}]$.
@@
$-\int_{x_{i-\frac{1}{2}}}^{x_{i+\frac{1}{2}}}\left(\boldsymbol{v}\left(x,y_{j+\frac{1}{2}}^{-}\right)\cdot\widehat{\boldsymbol{F}_{2}^{Q}}_{,j+\frac{1}{2}}\left(x,t\right)-\boldsymbol{v}\left(x,y_{j-\frac{1}{2}}^{+}\right)\cdot\widehat{\boldsymbol{F}_{2}^{Q}}_{,j-\frac{1}{2}}\left(x,t\right)\right)\mathrm{~{}d}x,$
@@
$\widehat{\boldsymbol{F}_{2}^{Q}}_{,j+\frac{1}{2}}\left(x,t\right):=\widehat{\boldsymbol{F}_{2}^{Q}}\left(\boldsymbol{U}_{h}\left(x,y_{j+\frac{1}{2}}^{-},t\right),\boldsymbol{U}_{h}\left(x,y_{j+\frac{1}{2}}^{+},t\right)\right),$
@@ 2.4 Fully discrete OEDG schemes
The combination of semi-discrete schemes (9) and (12) can be regarded as a system of ordinary differential equations (ODEs) with respect to time $t$. For brevity, we rewrite equations (9) and (12) in the following ODE form:
@@
where $\boldsymbol{U}_{h}$ denotes the combination of $\boldsymbol{R}_{h}$ and $\boldsymbol{Q}_{h}$. This ODE system (13) is typically discretized in time by a Runge–Kutta method, for example, the explicit third-order strong-stability-preserving (SSP) Runge–Kutta method:
@@
where $\boldsymbol{U}_{\sigma}^{n}$ is the OE numerical solution at the $n$-th time level, and we define $\boldsymbol{U}_{\sigma}^{0}=\boldsymbol{U}_{h}^{0}$ or $\boldsymbol{U}_{\sigma}^{0}=\mathcal{F}_{\text{OE}}\boldsymbol{U}_{h}^{0}$.
@@ 2.4.1 Component-wise OE procedure for $\boldsymbol{R}_{h}$
where $\tau$ is a pseudo-time different from $t$, ``"``"`$\circ$"` ` ∘ " denotes the Hadamard product of two vectors, $P^{m}$ is the standard $L^{2}$-projection operator into $\mathbb{V}_{R}^{m}$ for $m\geq 0$, and we define $P^{-1}=P^{0}$. The vector $\boldsymbol{\delta}_{ij}^{m}\left(\boldsymbol{R}_{h}^{n,\ell}\right)$ is defined by
@@
where $\beta_{ij}^{x}$ and $\beta_{ij}^{y}$ denote suitable estimates of the local maximum wave speed in the $x$- and $y$-directions, respectively, and $\boldsymbol{\sigma}_{i+\frac{1}{2},j}^{m}\left(\boldsymbol{R}_{h}\right)$ and $\boldsymbol{\sigma}_{i,j+\frac{1}{2}}^{m}\left(\boldsymbol{R}_{h}\right)$ correspond to the damping coefficients on the $x=x_{i+\frac{1}{2}}$ and $y=y_{j+\frac{1}{2}}$ edges, respectively. In practice, we take $\beta_{ij}^{x}$ and $\beta_{ij}^{y}$ as the spectral radii of the Jacobian matrices $\frac{\partial\boldsymbol{F}_{1}}{\partial{\boldsymbol{U}}}\left(\bar{\boldsymbol{U}}_{ij}^{n,\ell}\right)$ and $\frac{\partial\boldsymbol{F}_{2}}{\partial{\boldsymbol{U}}}\left(\bar{\boldsymbol{U}}_{ij}^{n,\ell}\right)$, respectively, and $\bar{\boldsymbol{U}}_{ij}^{n,\ell}$ denotes the cell average of $\boldsymbol{U}_{h}^{n,\ell}\left(\boldsymbol{x}\right)$ over $I_{ij}$. In equation (17), the damping coefficients are given by
$\boldsymbol{\sigma}_{i+\frac{1}{2},j}^{m}\left(\boldsymbol{R}_{h}\right)=\left(\sigma_{i+\frac{1}{2},j}^{m}\left(R_{h}^{\left(1\right)}\right),\sigma_{i+\frac{1}{2},j}^{m}\left(R_{h}^{\left(2\right)}\right),\cdots,\sigma_{i+\frac{1}{2},j}^{m}\left(R_{h}^{(6)}\right)\right)^{\top}$
$\boldsymbol{\sigma}_{i,j+\frac{1}{2}}^{m}\left(\boldsymbol{R}_{h}\right)=\left(\sigma_{i,j+\frac{1}{2}}^{m}\left(R_{h}^{\left(1\right)}\right),\sigma_{i,j+\frac{1}{2}}^{m}\left(R_{h}^{\left(2\right)}\right),\cdots,\sigma_{i,j+\frac{1}{2}}^{m}\left(R_{h}^{\left(6\right)}\right)\right)^{\top},$
@@
Note that the coefficient $\boldsymbol{\delta}_{ij}^{m}\left(\boldsymbol{R}_{h}^{n,\ell}\right)$ in the damping equations (16) only depends on the “initial” value $\boldsymbol{R}_{\sigma}\left(\boldsymbol{x},0\right)=\boldsymbol{R}_{h}^{n,\ell}\left(\boldsymbol{x}\right)$. Consequently, the damping equations (16) are actually a linear ODE system, and its exact solution can be analytically given. We now derive the exact solver of damping equations (16) for $\boldsymbol{R}_{\sigma}\left(\boldsymbol{x},\tau\right)$. Based on the local orthogonal basis $\left\{\phi_{ij}^{\left(\boldsymbol{\alpha}\right)}\left(\boldsymbol{x}\right):\left|\boldsymbol{\alpha}\right|\leq k\right\}$ of $\mathbb{P}^{k}\left(I_{ij}\right)$, the “initial” value $\boldsymbol{R}_{h}^{n,\ell}\left(\boldsymbol{x}\right)$ of the damping equations (16) can be expanded as
@@
Substitute (20)-(21) into (16) and take $v=\phi_{ij}^{\left(\boldsymbol{\alpha}\right)}\left(\boldsymbol{x}\right)$. For $\left|\boldsymbol{\alpha}\right|=\mu\geq 1$, one has
@@
Integrating (22) from $\tau=0$ to $\Delta t_{n}$ gives
@@
$\int_{I_{ij}}\left(\boldsymbol{R}_{\sigma}-P^{m-1}\boldsymbol{R}_{\sigma}\right)\phi_{ij}^{\left(\boldsymbol{\alpha}\right)}\left(\boldsymbol{x}\right)\mathrm{~{}d}\boldsymbol{x}=\boldsymbol{0}\quad\forall m\geq 0,$
@@
Remark 1.
In essence, the OE procedure (25) suppresses potential spurious oscillations in the DG solution by damping its modal coefficients. Since the OE procedure does not interfere with DG spatial discretization or RK stage updates, it can be easily incorporated into existing DG codes as an independent module and easily extensible onto general meshes. Thanks to the exact solver of the OE procedure, its implementation is straightforward and efficient. Furthermore, the OEDG method remains stable under the normal CFL condition, and many desirable properties of the original DG method are preserved, including conservation and optimal convergence rates (see the theoretical justification in [20] and the numerical evidence in section 4). Note that the OE procedure is applied directly to the conservative variables without the need for characteristic decomposition, which greatly reduces computational costs. As with the OE technique in [20], the present OE procedure is also scale-invariant and evolution-invariant, thereby being capable of effectively eliminating spurious oscillations for problems across different scales and wave speeds.
@@ 2.4.2 LDF OE procedure for $\boldsymbol{Q}_{h}$
$\begin{split}\boldsymbol{Q}_{\sigma}^{n,\ell}\left(\boldsymbol{x}\right)&=\left(\mathcal{F}_{\text{OE}}^{Q}\boldsymbol{Q}_{h}^{n,\ell}\right)\left(\boldsymbol{x}\right)=\boldsymbol{Q}_{\sigma}\left(\boldsymbol{x},\Delta t_{n}\right)\\ &=\sum_{\eta=0}^{D_{Q}^{0}-1}Q_{ij}^{\left(\eta\right)}\left(0\right)\boldsymbol{\varphi}_{ij}^{\left(\eta\right)}\left(\boldsymbol{x}\right)+\sum_{\mu=1}^{k}\exp\left(-\Delta t_{n}\sum_{m=0}^{\mu}\delta_{ij}^{m}\left(\boldsymbol{Q}_{h}^{n,\ell}\right)\right)\sum_{\eta=D_{Q}^{\mu-1}}^{D_{Q}^{\mu}-1}Q_{ij}^{\left(\eta\right)}\left(0\right)\boldsymbol{\varphi}_{ij}^{\left(\eta\right)}\left(\boldsymbol{x}\right).\end{split}$ (35)
Remark 2.
The OE procedure (35) not only effectively suppresses potential spurious oscillations but also naturally maintains the LDF property of the magnetic field. Moreover, this LDF OE procedure inherits many notable advantages of the component-wise OE procedure (25) mentioned in Remark 1, including: being non-intrusive, which enables easy integration into existing (LDF) DG codes; straightforward extensibility to general meshes; stability under normal CFL conditions; scale invariance and evolution invariance, which are crucial for effectively eliminating spurious oscillations across various scales and wave speeds; being free of characteristic decomposition; retaining the conservation and optimal convergence rates of the original DG method. All these advantages underscore the strengths of the LDF OE procedure.